At what interest rate must $5600 be compounded annually to equal $8687.44 after 9 yr? (Round to the nearest percent.)


Answer by stanbon(75887)

At what interest rate must $5600 be compounded annually to equal $8687.44 after 9 yr? (Round to the nearest percent.)
A=P(1+r/n)^(nt)
8687.44 = 5600(1+r/1)^(1(9))
1.55132857...=(1+r)^9
Take the 9th root or use logarithms to solve.
1.050000003...=1+r
r=0.05 or 5%
Cheers,
Stan H.